Below are the most recent balance sheets for Country Kettles, Inc. Excluding accumulated depreciation, determine whether each item is a source or a use of cash, and the amount (Input all amounts as positive values):

COUNTRY KETTLES, INC.
Balance Sheets
2016 2017
  Assets
  Cash $ 31,100 $ 30,260
  Accounts receivable 70,600 73,720
  Inventories 61,500 63,750
  Property, plant, and equipment 154,000 164,200
    Less: Accumulated depreciation 46,480 50,600
  Total assets $ 270,720 $ 281,330
  Liabilities and Equity
  Accounts payable $ 45,600 $ 47,760
  Accrued expenses 6,980 6,180
  Long-term debt 26,300 29,050
  Common stock 23,000 27,700
  Accumulated retained earnings 168,840 170,640
  Total liabilities and equity $ 270,720 $ 281,330
Answer #1
Liability= if increase Marked as Source, Decrease marked as Use
Asset = if increase Marked as Use, Decrease marked as Source
2016 2017
Difference(2016-2017) Source/Use
  Cash 31,100 30,260 840 Source Decrease
  Accounts receivable 70,600 73,720 -3,120 Use Increase
  Inventories 61,500 63,750 -2,250 Use Increase
  Property, plant, and equipment 107,520 113,600 -6,080 Use Increase
  Accounts payable 45,600 47,760 -2,160 Source Increase
  Accrued expenses 6,980 6,180 800 Use Decrease
  Long-term debt 26,300 29,050 -2,750 Source Increase
  Common stock 23,000 27,700 -4,700 Source Increase
  Accumulated retained earnings 168,840 170,640 -1,800 Source Increase
  Property, plant, and equipment= value of asset - Accumulated depreciation
  Property, plant, and equipment 2016= 154000-46480 =107520
  Property, plant, and equipment 2017= 164200-50600 =113,600